OpenStudy (anonymous):

in a right triangle find the length of the side not given b=2 c=sqrt29

OpenStudy (anonymous):



OpenStudy (anonymous):

c^2 = a^2 + b^2 c^2 - b^2 = a^2 29 - 4 = a^2 25 = a^2 a =5